How many liters of a 40%-alcohol solution must be mixed with 10 liters of a solution that is 80% alcohol to get a solution that is 60% alcohol?
Answer by RAY100(1637)

Let x = additive volume in liters
.
.40*x + .80*10 = .60*(10+x)
.
.4x +8 = 6 + .6x
.
2=.2x
.
10=x
.
check
.4*10 +.8*10 = .6 *(10+10)=12,,,,ok
.